OpenStudy (anonymous):

What is the value of 6x8-4[7+(8-3)] ? WILL GIVE MEDAL!!PLZ

OpenStudy (tkhunny):

What's the trick? Show us your best work with the Order of Operations.

OpenStudy (anonymous):

[7+(8-3)] 8-3=5 5+7+12

OpenStudy (anonymous):

\[6\times8-4[7+(8-3)] \\ 6\times 8-4[7+5] \] is the first step

OpenStudy (anonymous):

then as you had \[6\times 8-4\times 12\]

OpenStudy (anonymous):

then do both multiplications, finally subtract

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

yes thankyou!!